OpenStudy (anonymous):

O.k. this is a easy one,however,I've forgotten the process.Pls.refresh my memory on this equation: y-2+3y=10

OpenStudy (anonymous):

-2+4y=10 4y=12 y=3

OpenStudy (anonymous):

Pretend y is x.

OpenStudy (anonymous):

It's just a different variable.

OpenStudy (anonymous):

combine like terms, add 2 to both sides to get the y alone, divide by 4

OpenStudy (anonymous):

Unless, however, the problem is parametric, in which case the equation is used to find the y coordinate. The x equation is treated the same way.

OpenStudy (anonymous):

Can you type out the whole process for me to see?

OpenStudy (anonymous):

Solving for y?

OpenStudy (anonymous):

I believe so.

OpenStudy (anonymous):

\[Y-2+3Y=10\] \[Y+3y=10-2=8\] \[4y=8\] \[Y=8div4\] \[Y=2\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

I have to look it over before I can say whether or not it makes sense.

OpenStudy (anonymous):

So the Y is equal to 1 that made it 4?

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

No, it equals 2.

OpenStudy (anonymous):

In order for the equation to hold true.

OpenStudy (anonymous):

No,what I meant was : Y+ 3y = 4?

OpenStudy (anonymous):

No, y+3y=8 because you subtract 2 from 10 first, getting 8. So, y+3y=8 Which then becomes 4y because you add the y's together, or combine like terms. Then you get y=2, because you divide 8 by 4.

OpenStudy (anonymous):

O.k.,thanks for the clarification.

OpenStudy (anonymous):

You all set?

OpenStudy (anonymous):

Yeah,for now anyways.:)

OpenStudy (anonymous):

Haha cool. Glad to help.